# Observations of microlensed images with dual-field interferometry: on-sky demonstration and prospects

### v1

@@ 5.2 Luminous-lens Model
We found that the closure phase data do not provide strong evidence for the light from the lens. Including the lens light in the fits improves the $\chi^{2}$ by 1.3, 0.1, and 0.9 for epochs 1a, 1b, and 2, respectively. The corresponding 95% upper limits on $\eta_{\rm b}$ are 0.74, 0.81, and 0.44, respectively. Moreover, we noticed that the lens flux is correlated with the projected lens–source separation (and so, the flux ratio between the microlensed images) and the angular Einstein radius.
Thus, we explored the possibility that the light from the lens may be a source of tensions between the photometric and interferometric data discussed above. We repeated modeling, taking into account priors on $u$ from the light curve model. The best-fit parameters are reported in Table 7. While the tension between $u$ determined from the light curve model and that from the closure phase model is removed, the position angles are still slightly different.
Table 7: Best-fit Parameters of the Luminous-Lens Closure Phase Model
@@
$\chi^{2}$/d.o.f. $824.4/924$ $49.6/48$ $282.6/308$
5.3 Luminous-blend Model
@@
We also consider the model with a luminous blend, which has two additional parameters compared to the luminous-lens model, the offset of the blend $(\Delta\alpha_{\rm b},\Delta\delta_{\rm b})$ relative to the lens in the sky. We searched for possible blends on a grid of $201\times 201$ positions uniformly spread over the range $-20\leq\Delta\alpha_{\rm b},\Delta\delta_{\rm b}\leq 20\,\mathrm{mas}$. We kept the position of the blend fixed but allowed the other parameters ($u$, PA, $\theta_{\rm E}$, and $\eta_{\rm b}$) to vary. The best-fit parameters were found using a downhill approach using the Nelder–Nead algorithm. We adopted a prior on the lens–source separation $u$ from the light curve model, and we required $\eta_{\rm b}\geq 0$.
@@
The grid search results for the epoch 2 data are presented in Figure 7. There is only one local minimum around $(\Delta\alpha_{\rm b},\Delta\delta_{\rm b})=(-6.0,15.8)\,\mathrm{mas}$ that may be statistically significant (the $\chi^{2}$ statistics is improved by $\Delta\chi^{2}=26.6$). We explored this local minimum using the MCMC approach and found $u=0.4295\pm 0.0016$, $\mathrm{PA}=26.00\pm 0.24$ deg, $\theta_{\rm E}=1.2672\pm 0.0054$ mas, $\Delta\alpha_{\rm b}=-5.91\pm 0.18$ mas, $\Delta\delta_{\rm b}=15.73\pm 0.32$ mas, and $\eta_{\rm b}=0.0426\pm 0.0094$. Thus, even if we consider this solution statistically significant, the blend does not explain the tension between the position angle measured using the light curve and interferometric data.
@@
Figure 7: Detection map for possible luminous blends in the VLTI epoch 2 data. The color codes the $\chi^{2}$ improvement over the model without the blend. The lens is located in the origin of the coordinate system.
5.4 Final Closure Phase Models
@@
Finally, we measure the angular Einstein radius of the event $\theta_{\rm E}$ using all VLTI data combined. The baseline model has five parameters: the source–lens separation $u$ and position angle of the images PA during epochs 1 and 2, and $\theta_{\rm E}$ (the same during both epochs). To estimate the impact of systematic errors on the parameter uncertainties, we employ the bootstrapping method (e.g. Efron, 1979; Kervella et al., 2004; Lachaume et al., 2019). It involves randomly selecting eight VLTI interferograms (with replacement), feeding them to our fitting routine and then repeating the procedure multiple times (in this case, 5000 times) to obtain the multivariate probability density function for the parameters of the model. Such a procedure enables us to retrieve more realistic uncertainties on the model parameters. The results of the bootstrapping method are reported in the second column of Table 8.
@@
We then repeat the modeling, allowing an additional flux from the lens. We consider two models: with and without the prior on $u$ from the light curve model (third and fourth columns of Table 8.) As in Section 5.2, we find that including the lens light does not significantly improve the fits, and we can only measure upper limits on $\eta_{\rm b}$ (95% confidence).
Table 8: Final Closure Phase Models
@@ 6 Physical Parameters of the Lens
where $D_{\rm s}\approx 8\,\mathrm{kpc}$ is the source distance. As the interferometric data do not provide strong evidence for the light from the lens, we adopt $\theta_{\rm E}=1.2802\pm 0.0085$ mas (Table 8) as our final measurement. Using $\pi_{\rm E}=0.3330\pm 0.0084$, we find $M=0.472\pm 0.012\,M_{\odot}$, $\pi_{\rm rel}=0.426\pm 0.011$ mas, and $D_{\rm l}=1.81\pm 0.05$ kpc. The lens is, therefore, most likely a main-sequence star located in the nearby Galactic disk. Alternatively, it may be a white dwarf, but because white dwarfs are much less frequent than main-sequence stars, this possibility is unlikely.
If the lens is a main-sequence star, its absolute magnitude is $M_{K}=5.76\pm 0.04$ (Pecaut & Mamajek, 2013), and its apparent $K$-band magnitude is $17.20\pm 0.07$, assuming the $K$-band extinction of 0.15 mag toward the lens (Nataf et al., 2013). That corresponds to $\eta_{\rm b}\approx 0.03$ in the $K$-band, consistent with the limits on the lens flux from the VLTI data.
The trajectory of the source in the sky is presented in Figure 8 by a gray solid line. Open symbols (red square and blue circle) mark the position of the source during the two epochs of VLTI observations, whereas the filled symbols mark the positions of the images of the source.
Figure 8: Geometry of the event. The gray solid line marks the trajectory of the source in the sky. Open symbols (red square and blue circle) mark the position of the source during the two epochs of VLTI observations, whereas the filled symbols mark the positions of the images of the source. The dashed circle shows the Einstein radius.
@@ 7 Discussion and Conclusions
The start of the scientific operations of GRAVITY Wide opens up completely new possibilities for the follow-up of microlensing events. Only very few of the brightest events could have been observed with the standard on-axis GRAVITY mode. In contrast, GRAVITY Wide observations provide the opportunity for studies of a large sample of events and measurements of masses, distances, and transverse velocities of isolated objects, including neutron stars and black holes. Interferometric observations are very efficient at the same time—only one-hour exposure is sufficient for determining the angular Einstein radius. Thus, interferometric observations with GRAVITY Wide (and, its successor, GRAVITY+) hold promise for routine detections of isolated neutron stars and stellar remnants.
@@
In this paper, we report the first successful observation of a microlensing event with GRAVITY Wide and the resolution of microlensed images in the event OGLE-2023-BLG-0061/KMT-2023-BLG-0496. These observations serve as a testbed for verifying the capabilities of the new instrument. In particular, some observations were carried out using the low-resolution mode, which had never been tried before with GRAVITY Wide. The comparison of medium- and low-resolution interferograms (epochs 1a and 1b) reveals low-level systematic differences between the inferred parameters of the lensing system (Table 5). Similar systematic differences can be seen in models fitted to individual exposures taken during epoch 2 (Table 6). Judging from Figures 4 and 5, the systematic and correlated errors in the closure phase data can reach up to 5–10 deg. We defer a detailed investigation of systematic errors in the closure phase data to a separate study.
@@
We use the bootstrapping method to deal with these systematic and correlated errors. Thanks to a large number of observations, we are still able to measure the angular Einstein radius with a sub percent precision, $\theta_{\rm E}=1.2802\pm 0.0085$ mas. By combining the information from the light curve and the closure phase data, we measure the mass of the lens with a precision of 2.6%, $M=0.472\pm 0.012\,M_{\odot}$.
@@
Independent tests of the accuracy of the closure phase model are possible with the light curve model. The model of the microlensing event light curve predicts the brightness ratio of the microlensed images and their position angle in the sky. While the predicted flux ratio matches that inferred from the closure phase data reasonably well (Figure 6), there is a tension in the position angles measured using the epoch 2 data. This tension cannot be explained by additional blended light in the closure phase data (Section 5.3).
@@
One possible explanation for this systematic difference is the unaccounted systematic errors in the closure phase data. Alternatively, the problem may lie in the light curve model. However, we find this unlikely, because the residuals from the best-fit light curve model are smaller than $\approx 0.01$ mag, and there is no strong evidence that the light curve model is inadequate. A binary-lens model can be ruled out because the binary lens would produce additional images of the source, which should have been detected in our grid search. Still, the unaccounted orbital motion of the source (“xallarap effect”) could, in principle, partly explain the discrepancy. The xallarap effect would not create additional images but would deflect the path of the source in the sky. Our modeling, however, do not provide strong evidence for xallarap in the light curve data (Appendix A).
@@ Acknowledgements
Based on observations collected at the European Southern Observatory under ESO program 108.220D. This research was funded in part by National Science Centre, Poland, grants OPUS 2021/41/B/ST9/00252 and SONATA 2023/51/D/ST9/00187 awarded to P.M. This research is partly supported by the National Natural Science Foundation of China (Grant No. 12133005) and the science research grants from the China Manned Space Project with No. CMS-CSST-2021-B12. S.D. acknowledges the New Cornerstone Science Foundation through the XPLORER PRIZE.
GRAVITY+ is developed by the Max Planck Institute for extraterrestrial Physics, the Institute National des Sciences de l’Univers du CNRS (INSU) with its institutes LESIA / Paris Observatory-PSL, IPAG / Grenoble Observatory, Lagrange / Côte d’Azur Observatory and CRAL / Lyon Observatory, the Max Planck Institute for Astronomy, the University of Cologne, the CENTRA - Centro de Astrofisica e Gravitação, the University of Southampton, the Katholieke Universiteit Leuven and the European Southern Observatory. D.D. has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement CoG - 866070).
This research has made use of the KMTNet system operated by the Korea Astronomy and Space Science Institute (KASI) at three host sites of CTIO in Chile, SAAO in South Africa, and SSO in Australia. Data transfer from the host site to KASI was supported by the Korea Research Environment Open NETwork (KREONET). This research was supported by KASI under the R&D program (project No. 2024-1-832-01) supervised by the Ministry of Science and ICT. W.Zang, H.Y., S.M., R.K., J.Z., and W.Zhu acknowledge support by the National Natural Science Foundation of China (Grant No. 12133005). W.Zang acknowledges the support from the Harvard-Smithsonian Center for Astrophysics through the CfA Fellowship. J.C.Y. and I.-G.S. acknowledge support from U.S. NSF Grant No. AST-2108414. Y.S. acknowledges support from BSF Grant No. 2020740. Work by C.H. was supported by the grants of National Research Foundation of Korea (2019R1A2C2085965 and 2020R1A4A2002885). J.C.Y. acknowledges support from a Scholarly Studies grant from the Smithsonian Institution.
@@ Appendix A Xallarap Models
In this section, we explore whether the orbital motion of the source (xallarap effect) can explain the discrepancy between the position angle of the microlensed images measured using VLTI data and calculated from the light curve model. Compared to the standard point-source point-lens model, models including the xallarap effect have seven additional parameters describing the shape of the source’s orbit. These are $P_{\xi}$ – the orbital period, $a_{\xi}$ – the semimajor axis (expressed in Einstein radius units), $i_{\xi}$ – the inclination of the orbit, $\Omega_{\xi}$ – the longitude of the ascending node of the orbit, $u_{\xi}$ – the argument of latitude at the reference epoch (which is taken to be $t_{0,\xi}=t_{0,\rm par}$), $e_{\xi}$ – the eccentricity of the orbit, and $\omega_{\xi}$ – the argument of periapsis of the orbit.

### v2

@@ 5.2 Luminous Lens Model
We find that the closure-phase data do not provide strong evidence for the light from the lens. Including the lens light in the fits improves the $\chi^{2}$ by 1.3, 0.1, and 0.9 for epochs 1a, 1b, and 2, respectively. The corresponding 95% upper limits on $\eta_{\rm b}$ are 0.74, 0.81, and 0.44, respectively. Moreover, we notice that the lens flux is correlated with the projected lens–source separation (and so the flux ratio between the microlensed images) and the angular Einstein radius.
Thus, we explore the possibility that the light from the lens may be a source of the tensions between the photometric and interferometric data discussed above. We repeat the modeling, taking into account the priors on $u$ from the light-curve model. The best-fit parameters are reported in Table 7. While the tension between $u$ determined from the light-curve model and that from the closure-phase model is removed, the position angles are still slightly different.
Table 7: Best-fit Parameters of the Luminous Lens Closure-phase Model
@@
$\chi^{2}$/dof $824.4/924$ $49.6/48$ $282.6/308$
5.3 Luminous Blend Model
@@
We also consider the model with a luminous blend, which has two additional parameters compared to the luminous lens model, the offset of the blend $(\Delta\alpha_{\rm b},\Delta\delta_{\rm b})$ relative to the lens in the sky. We searched for possible blends on a grid of $201\times 201$ positions uniformly spread over the range $-20\leq\Delta\alpha_{\rm b},\Delta\delta_{\rm b}\leq 20\,\mathrm{mas}$. We kept the position of the blend fixed but allowed the other parameters ($u$, PA, $\theta_{\rm E}$, and $\eta_{\rm b}$) to vary. The best-fit parameters were found using a downhill approach using the Nelder–Nead algorithm. We adopted a prior on the lens–source separation $u$ from the light-curve model, and we required $\eta_{\rm b}\geq 0$.
@@
The grid search results for the epoch 2 data are presented in Figure 8. There is only one local minimum around $(\Delta\alpha_{\rm b},\Delta\delta_{\rm b})=(-6.0,15.8)\,\mathrm{mas}$ that may be statistically significant (the $\chi^{2}$ statistics is improved by $\Delta\chi^{2}=26.6$). We explored this local minimum using the MCMC approach and found $u=0.4295\pm 0.0016$, $\mathrm{PA}=26.00\pm 0.24$ deg, $\theta_{\rm E}=1.2672\pm 0.0054$ mas, $\Delta\alpha_{\rm b}=-5.91\pm 0.18$ mas, $\Delta\delta_{\rm b}=15.73\pm 0.32$ mas, and $\eta_{\rm b}=0.0426\pm 0.0094$. Thus, even if we consider this solution statistically significant, the blend does not explain the tension between the position angle measured using the light-curve and interferometric data.
@@
Figure 8: Detection map for possible luminous blends in the VLTI epoch 2 data. The color codes the $\chi^{2}$ improvement over the model without the blend. The lens is located at the origin of the coordinate system.
5.4 Final Closure-phase Models
@@
Finally, we measure the angular Einstein radius of the event $\theta_{\rm E}$ using all VLTI data combined. The baseline model has five parameters: the source–lens separations $u$ and position angles of the images PA during epochs 1 and 2 and $\theta_{\rm E}$ (the same during both epochs). To estimate the impact of systematic errors on the parameter uncertainties, we employ the bootstrapping method (e.g., Efron, 1979; Kervella et al., 2004; Lachaume et al., 2019). This involves randomly selecting eight VLTI interferograms (with replacement), feeding them to our fitting routine, and then repeating the procedure multiple times (in this case, 5000 times), to obtain the multivariate probability density function for the parameters of the model. Such a procedure enables us to retrieve more realistic uncertainties on the model parameters. The results of the bootstrapping method are reported in the second column of Table 8.
@@
We then repeat the modeling, allowing an additional flux from the lens. We consider two models: with and without the prior on $u$ from the light-curve model (the third and fourth columns of Table 8.) As in Section 5.2, we find that including the lens light does not significantly improve the fits, and we can only measure upper limits on $\eta_{\rm b}$ (95% confidence).
Table 8: Final Closure-phase Models
@@ 6 Physical Parameters of the Lens
where $D_{\rm s}\approx 8\,\mathrm{kpc}$ is the source distance. As the interferometric data do not provide strong evidence for the light from the lens, we adopt $\theta_{\rm E}=1.2802\pm 0.0085$ mas (Table 8) as our final measurement. Using $\pi_{\rm E}=0.3330\pm 0.0084$, we find $M=0.472\pm 0.012\,M_{\odot}$, $\pi_{\rm rel}=0.426\pm 0.011$ mas, and $D_{\rm l}=1.81\pm 0.05$ kpc. The lens is, therefore, most likely a main-sequence star located in the nearby Galactic disk. Alternatively, it may be a white dwarf. The mass function of field white dwarfs peaks at approximately $0.6\,M_{\odot}$ and falls steeply toward lower masses (e.g., Tremblay et al., 2016; Torres et al., 2021; Cunningham et al., 2024). Because white dwarfs less massive than $0.5\,M_{\odot}$ are much less frequent than main-sequence stars of that mass, we find this possibility unlikely.
If the lens is a main-sequence star, its absolute magnitude is $M_{K}=5.76\pm 0.04$ (Pecaut & Mamajek, 2013), and its apparent $K$-band magnitude is $17.20\pm 0.07$, assuming $K$-band extinction of 0.15 mag toward the lens (Nataf et al., 2013). That corresponds to $\eta_{\rm b}\approx 0.03$ in the $K$ band, consistent with the limits on the lens flux from the VLTI data.
The trajectory of the source in the sky is presented in Figure 9 by a gray solid line. The open symbols (red square and blue circle) mark the positions of the source during the two epochs of VLTI observations, whereas the filled symbols mark the positions of the images of the source.
Figure 9: Geometry of the event. The gray solid line marks the trajectory of the source relative to the lens (black cross) in the sky. The open symbols (red square and blue circle) mark the positions of the source during the two epochs of VLTI observations. The filled symbols (red squares and blue circles) mark the positions of the two images of the source observed during epoch 1 and epoch 2, respectively. The dashed circle shows the Einstein radius.
@@ 7 Discussion and Conclusions
The start of the scientific operations of GRAVITY Wide opens up completely new possibilities for the follow-up of microlensing events. Only very few of the brightest events could have been observed with the standard on-axis GRAVITY mode. In contrast, GRAVITY Wide observations provide the opportunity for studies of a large sample of events and measurements of masses, distances, and transverse velocities of isolated objects, including neutron stars and black holes. Interferometric observations are very efficient at the same time—only a 1 hr exposure is sufficient for determining the angular Einstein radius. That is in stark contrast to astrometric microlensing measurements, which require time-consuming monitoring lasting several years (Sahu et al., 2022; Lam et al., 2022). Astrometric measurements are also more susceptible to systematic errors due to blending than interferometric observations (Mróz et al., 2022). Thus, as argued by Gould (2023), interferometric observations with GRAVITY Wide (and its successor, GRAVITY+) hold promise for routine detections of isolated neutron stars and stellar remnants.
@@
In this paper, we report the first successful observation of a microlensing event with GRAVITY Wide and the resolution of microlensed images in the event OGLE-2023-BLG-0061/KMT-2023-BLG-0496. These observations serve as a test bed for verifying the capabilities of the new instrument. In particular, some observations were carried out using the low-resolution mode, which had never been tried before with GRAVITY Wide. The comparison of medium- and low-resolution interferograms (epochs 1a and 1b) reveals low-level systematic differences between the inferred parameters of the lensing system (Table 5). Similar systematic differences can be seen in models fitted to individual exposures taken during epoch 2 (Table 6). Judging from Figures 5 and 6, the systematic and correlated errors in the closure phase data can reach up to 5–10 deg. We defer a detailed investigation of the systematic errors in the closure-phase data to a separate study.
@@
We use the bootstrapping method to deal with these systematic and correlated errors. Thanks to the large number of observations, we are still able to measure the angular Einstein radius with a subpercent precision, $\theta_{\rm E}=1.2802\pm 0.0085$ mas. By combining the information from the light-curve and the closure-phase data, we measure the mass of the lens with a precision of 2.6%, $M=0.472\pm 0.012\,M_{\odot}$.
@@
Independent tests of the accuracy of the closure-phase model are possible with the light-curve model. The model of the microlensing event light curve predicts the brightness ratio of the microlensed images and their position angle in the sky. While the predicted flux ratio matches that inferred from the closure-phase data reasonably well (Figure 7), there is a tension in the position angles measured using the epoch 2 data. This tension cannot be explained by additional blended light in the closure-phase data (Section 5.3).
@@
One possible explanation for this systematic difference is the unaccounted-for systematic errors in the closure-phase data. Alternatively, the problem may lie in the light-curve model. However, we find this unlikely, because the residuals from the best-fit light-curve model are smaller than $\approx 0.01$ mag, and there is no strong evidence that the light-curve model is inadequate. A binary-lens model can be ruled out, because the binary lens would produce additional images of the source, which should have been detected in our grid search. Still, the unaccounted-for orbital motion of the source (the “xallarap effect”) could, in principle, partly explain the discrepancy. The xallarap effect would not create additional images but would deflect the path of the source on the sky. Our modeling, however, does not provide strong evidence for xallarap in the light-curve data (Appendix A).
The additional testing of the model presented in this paper will be possible thanks to the data from the Gaia satellite (Gaia Collaboration et al., 2016). OGLE-2023-BLG-0061/KMT-2023-BLG-0496 was observed and alerted by the Gaia Photometric Science Alerts system (Hodgkin et al., 2021) as Gaia23ckg. Time-series astrometric data collected by Gaia, when made public later this decade, may reveal deflection due to astrometric microlensing effects and enable independent measurements of the angular Einstein radius.
@@ Acknowledgements
We thank David Bennett and Łukasz Wyrzykowski for their comments on the manuscript.
Based on observations collected at the European Southern Observatory under ESO program 108.220D. This research was funded in part by the National Science Centre, Poland, grants OPUS 2021/41/B/ST9/00252 and SONATA 2023/51/D/ST9/00187 awarded to P.M. This research is partly supported by the National Natural Science Foundation of China (grant No. 12133005) and the science research grants from the China Manned Space Project with No. CMS-CSST-2021-B12. S.D. acknowledges the New Cornerstone Science Foundation through the XPLORER PRIZE.
GRAVITY+ is developed by the Max Planck Institute for Extraterrestrial Physics, the Institute National des Sciences de l’Univers du CNRS (INSU), with its institutes LESIA / Paris Observatory-PSL, IPAG / Grenoble Observatory, Lagrange / Côte d’Azur Observatory, and CRAL / Lyon Observatory, the Max Planck Institute for Astronomy, the University of Cologne, the CENTRA—Centro de Astrofisica e Gravitação, the University of Southampton, the Katholieke Universiteit Leuven, and the European Southern Observatory. D.D. has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement CoG—866070).
This research has made use of the KMTNet system operated by the Korea Astronomy and Space Science Institute (KASI) at the three host sites of CTIO in Chile, SAAO in South Africa, and SSO in Australia. Data transfer from the host site to KASI was supported by the Korea Research Environment Open NETwork (KREONET). This research was supported by KASI under the R&D program (project No. 2024-1-832-01) supervised by the Ministry of Science and ICT. W.Zang, H.Y., S.M., R.K., J.Z., and W.Zhu acknowledge support by the National Natural Science Foundation of China (grant No. 12133005). W.Zang acknowledges the support from the Harvard–Smithsonian Center for Astrophysics through the CfA Fellowship. J.C.Y. and I.-G.S. acknowledge support from U.S. NSF grant No. AST-2108414. Y.S. acknowledges support from BSF grant No. 2020740. The work by C.H. was supported by grants from the National Research Foundation of Korea (2019R1A2C2085965 and 2020R1A4A2002885). J.C.Y. acknowledges support from a Scholarly Studies grant from the Smithsonian Institution.
@@ Appendix A Xallarap Models
In this section, we explore whether the orbital motion of the source (xallarap effect) can explain the discrepancy between the position angle of the microlensed images measured using VLTI data and calculated from the light-curve model. Compared to the standard point-source point-lens model, models including the xallarap effect have seven additional parameters describing the shape of the source’s orbit. These are $P_{\xi}$ – the orbital period; $a_{\xi}$ – the semimajor axis (expressed in Einstein radius units); $i_{\xi}$ – the inclination of the orbit; $\Omega_{\xi}$ – the longitude of the ascending node of the orbit; $u_{\xi}$ – the argument of latitude at the reference epoch (which is taken to be $t_{0,\xi}=t_{0,\rm par}$); $e_{\xi}$ – the eccentricity of the orbit; and $\omega_{\xi}$ – the argument of periapsis of the orbit.